OpenStudy (anonymous):

The perimeter of a rectangular lot is 32 meters. The length exceeds the width by 4 meters. Find the length of the lot. The possible answers are: 6 m 14 m 10 m 18 m

OpenStudy (anonymous):

l=w+4 w+w+2(w+4)=32 4w+8=32 4w=24 w=6

OpenStudy (anonymous):

lenght=10